verify the identity
1.sin x sec x
2. cos x csc x =cot x
: verify the identity
1.sin x sec x
2. cos x csc x =cot x

Answer by stanbon(18718)

verify the identity
1.sin x sec x = sinx*(1/cosx) = sin/cos = tanx
---------------------------
2. cos x csc x =cot x
cos*(1/sin) = cos/sin = tanx
================
Cheers,
Stan H.